OpenStudy (anonymous):

Solve Sqrt7 - x = 3 for x? Is the answer -2, 2, 4, or -4?

OpenStudy (jdoe0001):

\(\bf \sqrt{7}-x=3?\)

OpenStudy (anonymous):

I believe it is \[\sqrt{7-x}=3\]

OpenStudy (jdoe0001):

tis ambiguous

OpenStudy (anonymous):

Because this gives an answer there matches the 4 given

OpenStudy (anonymous):

In order to solve this, then first you should try to get rid of the square root. This gives you:\[7-x=3^2\] now you want to isolate x, try and do that :)

OpenStudy (anonymous):

So the answer would be 4?

OpenStudy (anonymous):

No, you have \[3^2\] on the left side, that is 3*3 = 9, so you get \[7-x=9\]

OpenStudy (anonymous):

I meant on the right side.

OpenStudy (anonymous):

But it is asking me to solve for x. And -2, 2 ,4,and -4 are the only optional answers.

OpenStudy (anonymous):

Yes, and you have\[7-x=9\] if you subtract 7 on both sides then you get \[-x=2\]now you have that:\[x=\]

OpenStudy (anonymous):

Thank you so much for your time and help!!!!

OpenStudy (anonymous):

You're welcome :)